OpenStudy (anonymous):

Show why, when using the sieve method to find all primes less than n, we need only strike out multiples of primes whose square is less than or equal to n.

OpenStudy (anonymous):

in simple english, if you have a number larger than the square root that is a factor of \(n\) the other factor must be smaller than the square root

OpenStudy (anonymous):

so far I have Suppose to the contrary that \[n=ab\], where\[a,b>\sqrt{n}\] Then, \[n = ab > \sqrt{n}\ * \sqrt{n}\ = n\]this is a contradiction. I kind of understand it but I don't know how to put it in terms.